OpenStudy (anonymous):

y=tanx/x2 what is d/dy

OpenStudy (anonymous):

are you sure its d/dy and not d/dx?

OpenStudy (anonymous):

I'm assuming this is an implicit problem: \[\frac{d}{dy}(y=\frac{\tan(x)}{x^2})\] This gives: \[\frac{\sec(x)\tan(x)x^2-\tan(x)(2x)}{(x^2)^2}*\frac{dx}{dy}\] If it is d/dx. Then it is: \[\frac{dy}{dx}=\frac{\sec(x)\tan(x)x^2-\tan(x)(2x)}{(x^2)^2}\]

OpenStudy (anonymous):

here you can use the quotient rule or the product rule (tanx * (2x)^-1)